# RASP: A Drone-based Reconfigurable Actuation and Sensing Platform Towards Ambient Intelligent Systems

### v1

@@ 4.4. Drone Navigation
We use images from the camera network to guide and control the drone. We attach an aRuco marker to the top of the drone, just like the patterns used for landing the drone in Section 3. To move the drone to specific locations, we use the straight line path from the drone’s current location to the closest point of interest. Throughout our deployments in Section 5, we observed an median 2-D localization error of 3.29cm, using our camera network. Moreover, as shown in Figure 11, the localization error does not increase over time, as is common in inertial measurement unit and dead reckoning approaches.
Using additional markers on top of the drone and the straight line path to navigate is not ideal for generalizing to other drones and environments that are more winding. The focus of this portion of our work is to demonstrate the utility of RASP in an important application (personal assistant) and not specifically drone detection, navigation, or control. Hence, we leave these aspects to future work.
@@ 5. Deployment
We deployed the RASP based personal assistant into an office/lab setting as shown in Figure 11e. We used a network of 4 cameras in a uniform grid. As discussed in Section 4.1, the cameras self-localized themselves and generated the floor map according to (Liu et al., 2022) after an afternoon of observing people walking and moving through the space.
@@ 5.1. Benchmarking
Figure 13 highlights improvements that RASP brings compared to a purely static camera-based system across all sensing tasks, especially with respect to false positives. A successful or “accurate” trial in this context means that the system was able to correctly identify the correct object or location (object/location ID task), correctly identify the state of the object or location (object/location state task), or correctly identify when a targeted event occurs (surveillance task); any additional points of interest identified are counted as incorrect identifications (false positives). For the non-RASP solution, we run the entire SAM, LLaVA, DINO pipeline, but do not actuate the drone. Because we are often looking for small item(s) and locations in a large scene, the visual-language model pipeline often identifies multiple points of interest (e.g., DINO draws multiple bounding boxes and locations). Without a platform such as a drone that can “zoom in” and confirm, the sensing capabilities of this system is limited, and the false positive rate becomes extremely high and the precision becomes low with additional locations identified. However, adding in the RASP-equipped drone allows the system to actuate the drone to each location to obtain a closeup view of the location and remove extraneous locations or sense an aspect of the environment that a camera cannot (e.g., humidity). This both reduces false positives and improves overall accuracy. Table 2 breaks down the recall, precision, and f-1 score across all individual tasks to further illustrate improvements in true detection rate (recall) and precision by incorporating RASP.
@@ 5.2. Reconfiguring Mid-Mission
The previous section demonstrated how RASP could be used in conjunction with static sensors in the environment to better perform tasks requiring a single sensor or actuator in a home or office setting. Here, we look at scenarios where a user issues commands that require multiple sensing modalities and actuators, highlighting how RASP could be easily reconfigured mid-mission to satisfy these multi-layered and more complex tasks.
@@
Takeaways. We demonstrated how RASP could be used to reconfigure a drone, mid-mission, for tasks that require multiple sensors and/or actuators in a smart building setting with a decent success rate. However, a truly practical system still needs future improvements in drone landing (causes failed module swaps), methods for analyzing sensor data, and fine-grained control of drones.
@@ 6. Discussion
The focus of this work is to realize an autonomous reconfigurable and modular sensing and actuation platform drone platform that can adapt its sensing and actuation capabilities depending on the task at hand. As such, a single drone can be adapted for multiple tasks. We demonstrate the utility of such a platform for ambient intelligent systems by integrating it with a drone-based personal assistant that we envision will become commonplace in future homes, offices, and buildings. Specifically, we demonstrate the potential of leveraging static sensors in the environment and foundational models (e.g., visual-language models) in conjunction with a powerful and adaptable mobile platform in smart building, home, office, and indoor environments. To realize a fully autonomous drone-based personal assistant and other intelligent systems that can fully take advantage of RASP, there are several directions that need more exploration.
Robust path planning, navigation, and search for indoor environments. As discussed in Section 5.3, many of the tasks that failed were due to the simple search and navigation algorithms we implemented on the drone that were not able to capture and detect small objects of interest in a cluttered home/indoor setting. More work into exploration and detecting these small objects of interest is needed.
Penetrative AI, foundational models and leveraging diverse sensors in the environment. In this work, we leveraged cameras in conjunction with language to understand and interpret the environment. However, not all areas may have a camera (e.g., due to privacy or related concerns), and there are many more types of sensors that could be deployed; enabling foundational models and drones to understand the physical world (e.g., penetrative AI (Xu et al., 2024)) through all types of sensing modalities is important for generalizing our platform and case study to all types of scenarios.
Drone noise reduction, smaller sizes, and power efficiency. Drones and quadcopters are very loud because of its propellers. Having them comfortably co-exist with people in indoor environments requires more research into the design of them to reduce their size and noise, while prolonging their battery life and enabling automatic charging while docked.
@@ 7. Conclusion
We propose RASP, a modular and reconfigurable sensing and actuation platform that allows small drones to swap sensing and actuation capabilities dynamically. RASP allows a single drone to reconfigure itself to satisfy a diverse range of tasks that is commonly not possible with current drones. Current industrial drones are typically designed to satisfy a single application domain, and consumer drones are currently mostly used by hobbyists for entertainment. We demonstrate the diverse range of tasks RASP can enable by integrating RASP into a personal assistant system that also leverages static cameras to enable a wide range of useful tasks that were previously not possible. RASP is a step towards generalizing the use and presence of drones in daily life, much like how smartphones have become ubiquitous.
Acknowledgements.
This research was partially supported by COGNISENSE, one of seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA, as well as the National Science Foundation under Grant Number CNS-1943396. The views and conclusions contained here are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Columbia University, NSF, SRC, DARPA, or the U.S. Government or any of its agencies.

### v2

@@ 4.4. Drone Navigation
We use images from the camera network to guide the drone by attaching an aRuco marker to the top of the drone, just like the patterns used for landing the drone in Section 3. To move the drone to specific locations, we use the straight line path from the drone’s current location to the closest point of interest. Throughout our deployments in Section 5, we observed an median 2-D localization error of 3.29cm, using our camera network. Moreover, as shown in Figure 10, the localization error does not increase over time, as is common in inertial measurement unit and dead reckoning approaches. While there are more practical approaches for drone navigation, the focus of this work is to demonstrate the utility of RASP to LLMs that interact with the physical world. Hence, we leave these aspects to future work.
@@ 5. Deployment
We deployed our RASP based personal assistant into an office/lab setting as shown in Figure 10e, using a network of 4 cameras. As discussed in Section 4.1, the cameras self-localized themselves and generated the floor map according to (Liu et al., 2022) after an afternoon of observing people moving through the space. The goal of this deployment is to demonstrate improvements in task completion rate RASP provides to FMs for more general and less structured applications.
@@ 5.1. Benchmarking
Figure 12 highlights improvements that RASP brings compared to a purely static camera-based system across all sensing tasks, especially with respect to false positives. A successful or “accurate” trial in this context means that the system was able to correctly identify the correct object or location (object/location ID task), correctly identify the state of the object or location (object/location state task), or correctly identify when a targeted event occurs (surveillance task); any additional points of interest identified are counted as incorrect identifications (false positives). For the non-RASP solution, we run the entire SAM, LLaVA, DINO pipeline, but do not actuate the drone. Because we are often looking for small item(s) and locations in a large scene, the visual-language model pipeline often identifies multiple points of interest (e.g., DINO draws multiple bounding boxes and locations). Without a platform such as a drone that can “zoom in” and confirm, the sensing capabilities of this system is limited, and the false positive rate becomes extremely high and the precision becomes low with additional locations identified. However, adding in the RASP-equipped drone allows the system to actuate the drone to each location to obtain a closeup view of the location and remove extraneous locations or sense an aspect of the environment that a camera cannot (e.g., humidity). This both reduces false positives and improves overall accuracy. Table 2 breaks down the recall, precision, and f-1 score across all individual tasks to further illustrate improvements in true detection rate (recall). In total, integrating RASP with FMs improved the task success rate by $85\%$.
@@ 5.2. Reconfiguring Mid-Mission
The previous section demonstrated how RASP could be used in conjunction with static sensors in the environment to better perform tasks requiring a single sensor or actuator in a home or office setting. Here, we look at scenarios where a user issues commands that require multiple sensing modalities and actuators, highlighting how RASP could be easily reconfigured mid-mission to satisfy multi-layered tasks.
@@ 6. Discussion
The focus of this work is to demonstrate how a novel autonomous reconfigurable sensing and actuation drone platform is a critical step towards FMs, LLMs, and AGI that can interact, respond, and actuate our physical environments seamlessly, much like what has been achieved in the digital domain. To progress further, there are several directions that warrant further attention, which we discuss next.
Robust drone path planning, navigation, and search for indoor environments. As discussed in Section 5.3, many of the tasks that failed were due to the simple search and navigation algorithms we implemented that could not capture and detect small objects of interest in a cluttered home/indoor setting. More work into exploring, detecting, and navigating to these small objects of interest is needed.
Penetrative AI, foundation models and leveraging diverse sensors in the environment. In this work, we leveraged cameras in conjunction with language to understand and interpret the environment. However, not all areas may have a camera (e.g., due to privacy or related concerns), and there are many more types of sensors that could be deployed; enabling foundation models and drones to understand the physical world (e.g., penetrative AI (Xu et al., 2024)) through all types of sensing modalities is important for generalizing FMs and RASP to more scenarios.
@@ 7. Conclusion
We propose RASP, a modular and reconfigurable drone sensing and actuation platform that allows FMs and LLMs to dynamically to swap sensing and actuation capabilities on-the-fly. We demonstrate through a real deployment of a personal assistant application that RASP can improve the successful completion of complex tasks throughout our physical spaces by up to $85\%$. This drastic improvement stems from RASP’s ability to provide localized sensing and actuation in a large space even without large deployments of sensors and devices, as is commonly seen today. RASP is a critical step towards FMs and LLMs that can naturally interact and actuate the physical environment, just as they have shown in many applications in the digital domain.